# Frequency effects in Linear Discriminative Learning

### v1

@@ 9.1 Three methods for computing mappings in the DLM
WHL enables trial to trial learning and hence is, in principle, an excellent choice for datasets with learning events that are ordered in time. Examples of such datasets are child-directed speech in the CHILDES database ordered by the age of the addressee and the time-series of reaction times in mega-experiments (Heitmeier et al.,, 2022). The disadvantage of WHL is that it is computationally demanding, and prohibitively so for large datasets.
@@
Models using EL are not computationally demanding, but they are also not sensitive to the frequencies with which learning events occur. For usage-based approaches to language (see, e.g., Bybee,, 2010), this is a serious drawback. Nevertheless, EL has an important advantage of its own: it provides a window on what can be learned in principle, with infinite experience. In other words, EL is a good analytical tool for any datasets for which a type-based analysis is appropriate or insightful. For instance, if one's interest is in how well Dutch final devoicing can be mastered on the basis of subliminal learning only, the EL model informs us that a comprehension accuracy of 83% can be reached (see also Heitmeier et al.,, 2023). When measures are gleaned from an EL model and used as predictors for aspects of lexical processing, it will typically be necessary to include a measure of frequency of use as a covariate.
@@
An open question when it comes to comparing the three methods is whether measures of the models' accuracy should be type-based, i.e. all words contribute equally to the overall accuracy measure, or whether it should be token-based. Previous models have usually been evaluated based on type-accuracy, but a high token-accuracy in word recognition might be of more practical use in every day life than a high type- but low token-accuracy, where a model is able to recognise many low frequency words but struggles with many of the high frequency words with which speakers are constantly surrounded.
@@ 9.2 The relationship between word frequency and lexical decision reaction times
This finding fits in well with a recent debate that was re-sparked by Murray and Forster, (2004) arguing that rank-transformed frequencies account for lexical decision reaction times better than log-transformed ones and that, therefore, serial-search models should not be discounted as models of word recognition.888 However, note that the variance of word frequencies is similar in magnitude to the frequency itself (under the assumption that word frequencies are Poisson-distributed). By moving from frequencies to ranks, differences in frequency that seem large but that will vary substantially across samples are reduced to much smaller differences in ranks. In the light of these considerations, the excellent predictivity of rank for reaction times is due to distributional properties of the language in combination with sampling error, rather than due to serial searches in frequency ordered mental word lists. Recently, Kapatsinski, (2022) showed that log frequencies transformed by the logistic function (a function frequently used in deep learning models) predict reaction times in the same way as a rank-transformation, implying that the linear relationship between rank frequency and reaction times is not necessarily evidence in favour of the serial search model. Since the DLM is not a classification model, we do not make use of the logistic function but use correlation to compute how close a word's predicted meaning is to its true meaning. The estimated functional relation between -1000/RT and $1-r$ estimated with FIL is close to linear (see Figure 7, panel (2,1)), and a very similar partial effect emerges when the untransformed RTs are regressed on $1-r$. Thus, a linear relationship between a FIL-based predictor, $1-r$ (or equivalently, $r$) and reaction times falls out directly of the DLM, without requiring further transformations. This provides further evidence in favour of theories suggesting that frequency effects arise due to the distributional properties of words forms and meanings during learning.
@@
A question that we leave to future research is whether measures derived from FIL mappings will obviate the need to include frequency of use as a covariate, reduce the variable importance of this predictor, or complement it. One possible complication here is that while frequency-informed mappings make the DLM predictions more realistic, they naturally also create a confound with word frequency, which needs to be teased apart from other (e.g. semantic) effects captured by the DLM (see Gahl and Baayen,, 2023, for further discussion).
@@
What complicates such comparisons is that frequencies based on corpora typically concern lexical use in a language community, whereas individual speakers have very different experiences with lower-frequency words, depending on their interests, professions, education, and reading/speaking habits (e.g. Kuperman and Van Dyke,, 2013; Baayen et al.,, 2016). For instance, about 5% of the Dutch words in the present dataset are unknown to the fourth author. These are low-frequency words that the EL gets correct in 30 out of 35 cases. By contrast, the FIL gets only 2 out of these 35 cases correct. In a similar vein, Diependaele et al., (2012) showed that in the DLP, when reacting twice to the same lowest-frequency word stimuli in the DLP they only agree in about 50% of cases, bringing their accuracy to chance level.
@@
The inter- and intra-variability of subjects' word knowledge might also be a reason why FIL underestimates the variance of reaction times in the Dutch Lexicon Project for low-frequency words. Naturally, FIL, being based on `community' frequencies, is not able to account for this effect. Our results, however, highlight the importance of modelling not only the mean but also the variance of reaction times, as it might provide a further window into speakers' word recognition process and differentiating between different models attempting to account for the observed behavioural data.
@@
When evaluating the accuracy of a FIL model, having the closest proximity to the semantic vector of the target word is probably too stringent a criterion. Especially for lower-frequency words, having a rank among the top $k$ (for low $k$) nearest neighbors may be a more reasonable criterion. The reason is that for lower frequency words, knowledge of what words may mean is only approximate. For instance, although the names of gemstones such as jasper, ruby, tourmaline, and beryl will be known to many readers, picking out the ruby from four reddish pieces of rock requires more precise knowledge than is available to the authors. Especially for lexical decision making, being `close enough' may be good enough for a yes-response (see Heitmeier et al.,, 2022, for detailed discussion of lexical decision making).
@@ 9.3 Computational modelling of word frequency effects
Our findings are particularly enlightening when comparing them to accounts of word frequency effects in previous models of word recognition. To answer the three key questions regarding the word frequency effects identified in the introduction, in the DLM, frequency effects arise as a consequence of the input distribution it is trained on. Similarly to other network models, the effect of word frequency is stored in the model's mapping weights. Since the DLM is an extremely simple computational model involving only linear mappings, a fast closed-form solution is available, which, as demonstrated, is very similar to training the model incrementally on unordered data (as is done in all deep learning models utilising backpropagation, such as the triangle model Seidenberg and McClelland,, 1989). As such, it enables the modeller to answer many questions related to the word frequency effect, such as that of individual differences, in a computationally very lean way. Finally, we demonstrate that the DLM is able to account for the non-linear relationship between frequency and reaction times in lexical decision without requiring to use a log-transformation (cf. McClelland and Rumelhart,, 1981; Rumelhart and McClelland,, 1982; Seidenberg and McClelland,, 1989). Moreover, the DLM can not only account for frequency effects but also enables investigation of the influence of words' more finegrained characteristics on reaction times. Crucially, these are not independent. The DLM's predicted reaction times also depend on words' form and meaning properties, not only their frequency distribution.
@@
To put the present results in a broader perspective, consider the characterization given by Breiman et al., (2001) of machine learning on the one hand and statistics on the other hand. They argue that machine learning aims at obtaining accurate predictions. How these predictions are obtained and why a technique works is not of interest. By contrast, statistics aims to formulate a model that could have generated the data. Within this characterisation — which may be too extreme (e.g. Shmueli,, 2010) — FIL is much closer to statistical modeling than to machine learning, and it is surprising to us how much can be achieved simply with a form of weighted multivariate multiple regression. We hope that FIL will prove to be a useful tool not only for modeling data with `community' frequencies but also for exploring, by means of simulation, what the consequences are of individual differences in usage for lexical processing.

### v2

@@ 6.1 Three methods for computing mappings in the DLM
WHL enables trial to trial learning and hence is, in principle, an excellent choice for datasets with learning events that are ordered in time. Examples of such datasets are child-directed speech in the CHILDES database ordered by the age of the addressee and the time-series of reaction times in mega-experiments (Heitmeier et al., 2023a, ). The disadvantage of WHL is that it is computationally demanding, and prohibitively so for large datasets.
@@
Models using EL are not computationally demanding, but they are also not sensitive to the frequencies with which learning events occur. For usage-based approaches to language (see, e.g., Bybee,, 2010), this is a serious drawback. Nevertheless, EL has an important advantage of its own: it provides a window on what can be learned in principle, with infinite experience. In other words, EL is a good analytical tool for any datasets for which a type-based analysis is appropriate or insightful. For instance, if one's interest is in how well Dutch final devoicing can be mastered on the basis of subliminal learning only, the EL model informs us that a comprehension accuracy of 83% can be reached (see also Heitmeier et al., 2023b, ). When measures are gleaned from an EL model and used as predictors for aspects of lexical processing, it will typically be necessary to include a measure of frequency of use as a covariate.
@@
An open question when it comes to comparing the three methods is whether measures of the models' accuracy should be type-based, i.e. all words contribute equally to the overall accuracy measure, or whether it should be token-based. Previous models have usually been evaluated based on type-accuracy, but a high token-accuracy in word recognition might be of more practical use in every day life than a high type- but low token-accuracy, where a model is able to recognise many low frequency words but struggles with many of the high frequency words with which speakers are constantly confronted. In this context, it is worth keeping in mind the forgetting curve of Ebbinghaus (Ebbinghaus,, 1885) and recent advanced methods for repeating words at exactly the right moment in time to optimize fact learning, including vocabulary learning (van der Velde et al.,, 2022). From this literature, it is crystal clear that words encountered only once, which in corpora typically constitute roughly 50% of word types, cannot be learned with a single exposure. Whereas EL does not capture effects of frequency, FIL clarifies that once we start taking frequency into account, it is practice that makes perfect.
@@ 6.2 The relationship between word frequency and lexical decision reaction times
This finding fits in well with a recent debate that was re-sparked by Murray and Forster, (2004) arguing that rank-transformed frequencies account for lexical decision reaction times better than log-transformed ones and that, therefore, serial-search models should not be discounted as models of word recognition. [^9] However, note that the variance of word frequencies is similar in magnitude to the frequency itself (under the assumption that word frequencies are Poisson-distributed). By moving from frequencies to ranks, differences in frequency that seem large but that will vary substantially across samples are reduced to much smaller differences in ranks. In the light of these considerations, the excellent predictivity of rank for reaction times is due to distributional properties of the language in combination with sampling error, rather than due to serial searches in frequency ordered mental word lists. Recently, Kapatsinski, (2022) showed that log frequencies transformed by the logistic function (a function frequently used in deep learning models) predict reaction times in the same way as a rank-transformation, implying that the linear relationship between rank frequency and reaction times is not necessarily evidence in favour of the serial search model. Since the DLM is not a classification model, we do not make use of the logistic function but use correlation to compute how close a word's predicted meaning is to its true meaning. The estimated functional relation between -1000/RT and $1-r$ estimated with FIL is close to linear (see Figure 7, panel (2,1)), and a very similar partial effect emerges when the untransformed RTs are regressed on $1-r$. Thus, a linear relationship between a FIL-based predictor, $1-r$ (or equivalently, $r$) and reaction times falls out directly of the DLM, without requiring further transformations. This provides further evidence in favour of theories suggesting that frequency effects arise due to the distributional properties of words' forms and meanings during learning.  [^10] See Norris, (2006) for a similar argument in the context of a Bayesian approach to the role of word frequency in lexical processing.
@@
A question that we leave to future research is whether measures derived from FIL mappings will obviate the need to include frequency of use as a covariate, reduce the variable importance of this predictor, or complement it. One possible complication here is that while frequency-informed mappings make the DLM predictions more realistic, they naturally also create a confound with word frequency, which needs to be teased apart from other (e.g. semantic) effects captured by the DLM (see Gahl and Baayen,, 2023, for further discussion). Furthermore, in the present study we did not take into account (any measure of) a word's context (e.g. Adelman and Brown,, 2008; Hollis,, 2020). Gahl and Baayen, (2023) explore an additional measure of frequency which is informed by a word's context by measuring how strongly a word predicts itself in any utterances it occurs in. This measure outperforms frequency for predicting acoustic durations of English homophones, and it also outperforms frequency as a predictor of visual lexical decision latencies. This measure may complement the measures from the DLM.
@@
What complicates such comparisons is that frequencies based on corpora typically concern lexical use in a language community, whereas individual speakers have very different experiences with lower-frequency words, depending on their interests, professions, education, and reading/speaking habits (e.g. Kuperman and Van Dyke,, 2013; Baayen et al.,, 2016). For instance, about 5% of the Dutch words in the present dataset are unknown to the fourth author. These are low-frequency words that the EL gets correct in 30 out of 35 cases. By contrast, the FIL gets only 2 out of these 35 cases correct. In a similar vein, Diependaele et al., (2012) showed that in the DLP, when reacting twice to the same lowest-frequency word stimuli in the DLP participants only agree in about 50% of cases, bringing their accuracy to chance level.
@@
The inter- and intra-variability of subjects' word knowledge might also be a reason why FIL underestimates the variance of reaction times in the Dutch Lexicon Project for low-frequency words. Naturally, FIL, being based on `community' frequencies, is not able to account for this effect. Our results, however, highlight the importance of modelling not only the mean but also the variance of reaction times, as it might provide a further window into speakers' word recognition process and differentiating between different models attempting to account for the observed behavioural data. To our knowledge, previous models of word recognition have not attempted to account for both the mean and variance of reaction times (but see Ratcliff et al.,, 2004, for a model attempting to also predict the distribution of reaction times).
@@
When evaluating the accuracy of a FIL model, having the closest proximity to the semantic vector of the target word is probably too stringent a criterion. Especially for lower-frequency words, having a rank among the top $k$ (for low $k$) nearest neighbors may be a more reasonable criterion. The reason is that for lower frequency words, knowledge of what words may mean is only approximate. For instance, although the names of gemstones such as jasper, ruby, tourmaline, and beryl will be known to many readers, picking out the ruby from four reddish pieces of rock requires more precise knowledge than is available to the authors. Especially for lexical decision making, being `close enough' may be good enough for a yes-response (see Heitmeier et al., 2023a, , for detailed discussion of lexical decision making).
@@ 6.3 Computational modelling of word frequency effects
Our findings are particularly enlightening when comparing them to accounts of word frequency effects in previous models of word recognition. To answer the three key questions regarding the word frequency effects identified in the introduction, in the DLM, frequency effects arise as a consequence of the input distribution it is trained on. Similarly to other network models, the effect of word frequency is stored in the model's mapping weights. The DLM is an extremely simple computational model involving only linear mappings, which can be efficiently numerically computed. As demonstrated, the resulting mappings are very similar to those obtained by training the model incrementally on unordered data (as is done in all deep learning models utilising backpropagation, such as the triangle model, Seidenberg and McClelland,, 1989). As such, it enables the modeller to answer many questions related to the word frequency effect, such as that of individual differences, in a computationally lean way. Finally, we demonstrate that the DLM is able to account for the non-linear relationship between frequency and reaction times in lexical decision without requiring to use a log-transformation (cf. McClelland and Rumelhart,, 1981; Rumelhart and McClelland,, 1982; Seidenberg and McClelland,, 1989). Moreover, the DLM can not only account for frequency effects but also enables investigation of the influence of words' more finegrained characteristics on reaction times. Crucially, these are not independent. The DLM's predicted reaction times also depend on words' form and meaning properties, not only their frequency distribution.
Therefore, it is important to note that the DLM's performance also depends on many modelling choices, such as the chosen form granularity, semantic vectors etc. Ideal modelling choices can often vary across languages — for instance, while for English, Dutch and German, trigrams are often the unit of choice (Heitmeier et al., 2023a, ; Heitmeier et al.,, 2021), previous work has found that for Vietnamese, bigrams are preferable (Pham and Baayen,, 2015), while for Maltese, Kinyarwanda and Korean, form representations based on syllables perform well ( Nieder et al.,, 2023; van de Vijver et al.,, 2023, Chuang et al.,, 2022; an in-depth discussion of the various considerations when modelling a language with the DLM can be found in Heitmeier et al.,, 2021). While the present study is limited to Dutch, Mandarin and English, future work should further verify the efficacy of FIL on morphologically more diverse languages.
@@
To put the present results in a broader perspective, consider the characterization given by Breiman, (2001) of machine learning on the one hand and statistics on the other hand. They argue that machine learning aims at obtaining accurate predictions. How these predictions are obtained and why a technique works is not of interest. By contrast, statistics aims to formulate a model that could have generated the data. Within this characterisation — which may be too extreme (e.g. Shmueli,, 2010) — FIL is much closer to statistical modeling than to machine learning, and it is surprising to us how much can be achieved simply with a form of weighted multivariate multiple regression. We hope that FIL will prove to be a useful tool not only for modeling data with `community' frequencies but also for exploring, by means of simulation, what the consequences are of individual differences in usage for lexical processing.